If cosA= 3/5, find tanA as a fraction
Answer by Alan3354(69443)

If cosA= 3/5, find tanA as a fraction
-----------
sin^2 + cos^2 = 1 --> sin(A) = 4/5
---
tan = sin/cos
tan(A) = 4/3 (if in Q1)
tan(A) = -4/3 (if in Q4)